solve by factoring
4xsquared-12=22

Answer by dabanfield(803)

4x^2 - 12 = 22
4x^2 - 36 = 0
4*(x^2 - 9) = 0
4*(x-3)*(x+3) = 0
solutions are x-3 = 0 and x+3 = 0:
x=3 and x=-3